OpenStudy (anonymous):

The first stage of a rocket burns 28 seconds longer than the second stage. If the total burning time for both stages is 152 seconds then how long does each stage burn? So, how do you solve this?

OpenStudy (anonymous):

let x be the 1st stage y is the 2nd stage therefore x = 28 + y Since the total time is 152 sec, then x + y = 152 use 28 + y in place of x 28 + y + y = 152 solve for y that will be the burning time for stage 2 add 28 to that value to give you the burning time fr the 1st stage